# Optimal Rates for DP-SCO with a Single Epoch and Large Batches

### v1

@@ 1 Introduction
[10, 2, 30] have looked at various extensions to the $\ell_{2}$-DP-SCO problem. For example, DP-SCO guarantees under the loss functions being $\ell_{1}$-Lipschitz, or SCO properties of second-order-stationary points. These results, while being important for the DP-SCO literature, are not comparable to our line of work.
Practical enhancements, and empirical evaluation: Along with demonstrating the theoretical benefits of Accelerated-DP-SRGD , we demonstrate that the core ideas in our algorithm can be used to improve well-optimized benchmark results for DP learning [38, 23, 15]. While we chose $\eta_{t}=t$ in Accelerated-DP-SRGD to get optimal DP-SCO bounds, in practice we found setting $\eta_{t}=e^{\lambda t}$ and optimizing for $\lambda$ provides the best results. With this reparameterization, we make the crucial observation that if $\lambda$ is very high and $\tau_{t}=1$, then the Nesterov’s update simplifies to standard SGD update. Hence, there exists a parameterization under which Accelerated-DP-SRGD is equivalent to DP-FTRL  (differentially private follow-the-regularized-leader) [38, 23]. Hence, with more optimized parameters we can only get a better algorithm. The exact algorithm we use for empirical evaluation is given in Algorithm 3 (Algorithm DP-SRG-MEMF ).
In our empirical evaluation, we provide results for multi-class logistic regression on MNIST and CIFAR-10 as a proof-of-concept. We show that Accelerated-DP-SRGD provides a utility benefit over the prior SoTA [15] for the no-amplification [^7] Prior works [38, 43] mentioned that in distributed learning environments like federated learning, privacy amplification by sampling [36] is undesirable, and even when it is applicable, the true privacy guarantee depends on the precise implementation [16]. Hence, we focus on the evaluation of our algorithms when it does not have access to any form of privacy amplification. (i.e., streaming) setting that has been optimized for over a series of work [38, 23, 15, 14, 13]. A detailed empirical exploration of Accelerated-DP-SRGD and its practical variant on large-scale practical tasks is left for future exploration.
1.1 Our Contributions
@@
    Obtaining optimal DP-SCO errors with improved number of batch gradient steps (Theorem 4.6): We show that Accelerated-DP-SRGD can achieve optimal DP-SCO errors for smooth convex losses, in a single epoch with batch gradient steps of $T\approx\sqrt{n}$ (batch size $B=n/T)$, and exactly two gradient calls per data point. We required two technical ideas to achieve this: a) reducing the learning rate of non-private Nesterov’s method to account for the noise added due to privacy, and b) show that there exists appropriate choice of parameters for the algorithm under which the SRGs are low-sensitive (i.e., does not change too much by changing one data sample in the data set $D$).
  3. 3.
    Surpassing SoTA on benchmark tasks (Section 5): Inspired by our Accelerated-DP-SRGD , we create a heuristic (via exponential re-weighting of gradient differences) that empirically outperforms the highly optimized class of DP-FTRL algorithms [38, 23, 15] on multi-class logistic regression on MNIST and CIFAR-10.
@@ 2 Background on (Private) Learning
2.1 Definition for Single Epoch Training
Definition 2.1 (Single epoch).
A gradient based algorithm $\mathcal{A}(D)\to\{s_{1},\ldots,s_{T}\}$ is called single epoch, if it computes the gradients on any data point $d\in D$ at exactly one state $s_{t}$. Without additional restrictions, a state is allowed to encode arbitrary amount of history, i.e., $s_{t}\supseteq\{s_{1},\ldots,s_{t-1}\}$.
Notice that 2.1 does not include any restriction on the memory of the algorithm. While stating our results in the single epoch setting, we additionally incorporate the batch size, and the total number gradients evaluated per data point.
2.2 DP-Matrix Factorization (DP-MF ) and Continual Counting
Continual counting under DP [19, 24] is one of the most important primitives in the DP literature. Given a sequence of adaptively chosen data vectors $D=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{T}\}$ (with each $\left\|\mathbf{v}_{i}\right\|_{2}\leq 1$), the objective is to output $\mathbf{z}=\left[\sum\limits_{\tau\leq t}\mathbf{v}_{\tau}\right]_{t\in[T]}$ while preserving DP. It is well-known that one can output $\mathbf{z}_{\texttt{priv}}$ s.t. $\mathbb{E}[\|\mathbf{z}-\mathbf{z}_{\texttt{priv}}\|_{\infty}]=\text{polylog}(T)$. This can be achieved via the classic binary tree mechanism [19, 24, 32], see Lemma 4.3. We use the binary tree primitive for our theoretical guarantees. A recent line of work [23, 42, 26, 33, 34, 5] optimizes the error up to additive constant via factorization of a lower-triangular matrix of all ones $\mathbf{A}=\mathbf{B}\cdot\mathbf{C}$. For our empirical results, we appeal to the computational approach of [23] for finding the best factorizations ($\mathbf{B}$ and $\mathbf{C}$) w.r.t. the error in continual counting under DP.
@@ 2.4 DP-FTRL : DP follow the regularized leader
DP-FTRL  [49, 38, 23] is another standard approach for optimization, where in its simplest form, the state update is exactly that of DP-SGD when $\mathcal{C}=\mathbb{R}^{d}$, except the noise in two different rounds $\mathbf{b}_{t},\mathbf{b}_{t^{\prime}}$ are not independent. Instead (in the one-dimensional setting) $(\mathbf{b}_{1},\mathbf{b}_{2},\ldots\mathbf{b}_{T})\sim\mathcal{N}(0,\left(\mathbf{C}^{\top}\mathbf{C}\right)^{-1})$, where $\mathbf{C}$ is computed via DP-MF . While DP-FTRL has strong empirical properties comparable to DP-SGD , it is unknown whether it can achieve optimal DP-SCO error unless under additional restrictions (e.g., realizability [3]).
@@ 3 Our Algorithm
Our algorithm is a DP variant of Nesterov accelerated SGD [44, 9] using stochastic recursive gradients (SRGs): similarly to [45, 55] we maintain $\nabla_{t}$, an unbiased estimate of $\nabla F(\mathbf{x}_{t})$, and use $\nabla_{t}$ and batch $t+1$ to obtain $\nabla_{t+1}$. Our algorithm incorporates noise $\mathbf{b}_{t}$ for DP. We first give a utility analysis for a fixed value of the noise $\mathbf{b}_{t}$, and then in Section 4.2 we will specify how to set the noise to simultaneously achieve good privacy and utility. As mentioned in Section 1, we will use the DP-MF  machinery (Section 2.2) to set the $\mathbf{b}_{t}$’s.
It is worth mentioning that all three aspects of our algorithm are crucial: a) SRGs allow ensuring that the $\ell_{2}$-sensitivity of the gradient estimate at time $t$ is a factor of $1/t$, lower than that of the direct minibatch gradient ($\frac{1}{B}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t};d)$), b) Noise addition via DP-MF  allows the standard deviation of the noise $\mathbf{b}_{t}$ to grow only polylogarithmically in $T$, as opposed to $\sqrt{T}$ via simple composition analysis [25], and c) Accelerated SGD allows appropriately balancing the variance introduced by the SRG estimates, and that due to the i.i.d. sampling of the minibatches from the population. In fact in Section 4.3 we argue that it is necessary to have an accelerated SGD algorithm to achieve a single epoch optimal DP-SCO algorithm with batch size $B=\omega(1)$, if we were to use the SRGs as the gradient approximators.
@@
Inputs: Steps: $T$, batches: $\{\mathcal{B}_{t}\}_{0\leq t<T}$ of size $B$, learning rates: $\{\eta_{t}\}_{0\leq t<T}$, scaling $\beta$, interpolation params: $\{\tau_{t}\}_{t\in[T]}$, constraint set: $\mathcal{C}$ with $0\in\mathcal{C}$.
1:$\mathbf{x}_{0},\mathbf{z}_{0}\leftarrow 0$.
2:for $t\in\{0,\ldots,T-1\}$ do
3:     $\Delta_{t}\leftarrow\frac{1}{B}\sum_{d\in\mathcal{B}_{t}}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)]$
4:     $\nabla_{t}\leftarrow\frac{\sum_{i\leq t}\Delta_{i}}{\eta_{t}}$ (equivalently, $\nabla_{t}\leftarrow\frac{\Delta_{t}}{\eta_{t}}+\frac{\eta_{t-1}}{\eta_{t}}\cdot\nabla_{t-1}$)
5:     $\mathbf{z}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{z}_{t}-\frac{\eta_{t}}{\beta}\nabla_{t}+\mathbf{b}_{t}\right)$, where $\mathbf{b}_{t}$ is the noise added to ensure privacy.
@@
6:     $\mathbf{y}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}\right)$.
7:     $\mathbf{x}_{t+1}\leftarrow\left(1-\tau_{t+1}\right)\mathbf{y}_{t+1}+\tau_{t+1}\mathbf{z}_{t+1}$.
8:end for

### v2

@@ 1 Introduction
[9, 2, 25] have looked at various extensions to the $\ell_{2}$-DP-SCO problem. For example, DP-SCO guarantees under the loss functions being $\ell_{1}$-Lipschitz, or SCO properties of second-order-stationary points. These results, while being important for the DP-SCO literature, are not comparable to our work.
Future directions: While our work significantly improves on the SoTA in terms of gradient complexity and batch gradient steps in DP-SCO for smooth convex losses, it raises a lot of interesting open questions:
  1. (a)
    Can we obtain the same tighter batch gradient steps guarantee in Table 1, without assuming the unconstrained minimizer to be in the set $\mathcal{C}$? Informally, we only use this assumption to show the gradient norm is decreasing, i.e. the distance between $\mathbf{x}_{t}-\mathbf{x}_{t-1}$ is decreasing in $t$, and so the gradient differences $\Delta_{t}$ are bounded by smoothness. However, if gradients remain large during training, we expect training to reach the boundary of $\mathcal{C}$ quickly, in which case $\mathbf{x}_{t}-\mathbf{x}_{t-1}$ should also be small (because the gradient step is mostly reversed by the projection into $\mathcal{C}$). Hence we believe this assumption is not required for Accelerated-DP-SRGD to achieve $T=n^{1/4}$, but we do not know how to formally show this.
  2. (b)
    In our paper we only focused on $O(1)$-smooth convex losses. To the best of our knowledge, there is no single-epoch algorithm for the non-smooth setting. The best result we are aware of in this setting is that of [14], which uses $O(n)$ gradient oracle calls when e.g. $p,\varepsilon$ are constant, but can require up to $O(n^{4/3})$ gradient oracle calls for higher dimensions. It remains an open question what the smallest dimension-independent gradient complexity and batch gradient complexity required for optimal DP-SCO bounds in the non-smooth setting is. Our approach seems to require smoothness as the main primitive, stochastic recursive gradients, cares about differences of gradients at different points. So it is unlikely our techniques can be used to achieve similar improvements in the non-smooth setting.
  3. (c)
    Finally, while we believe the batch gradient steps bound of $n^{1/4}$ is tight for DP-SCO, we do not know how to prove a lower bound. Standard lower bounding techniques from the non-private literature [48] for batch gradient steps break down under DP setting as they operate in the regime where the dimensionality $p=\Omega(n^{2})$. In such high-dimensional regime, DP-SCO guarantees are vacuous. A well known result of [39] states that that no first-order method (i.e., algorithm that is only allowed to make gradient queries to points in the span of all gradients received so far) can achieve an empirical loss bound of $o(1/T^{2})$ in $T$ gradient queries for convex smooth losses, which even ignoring privacy suggests $T=n^{1/4}$ iterations are needed for any algorithm’s empirical loss to match the optimal stochastic error $1/\sqrt{n}$. It would be surprising if a DP algorithm could achieve optimal rates in $o(n^{1/4})$ iterations, as in the limit when $\varepsilon\rightarrow\infty$ it would beat this non-private lower bound. However, it is not obvious how to extend this lower bound to DP algorithms, as the noise addition violates the constraint that only points in the span of gradients are queried.
1.1 Our Contribution
@@
    Obtaining optimal DP-SCO errors with improved number of batch gradient steps (Theorem 4.2): We show that Accelerated-DP-SRGD can achieve optimal DP-SCO errors for $O(1)$-smooth convex losses, in a single epoch with batch gradient steps of $T\approx n^{1/4}$ (batch size $B=n/T$), if the unconstrained population risk minimizer $\mathbf{x}^{\dagger}$ is within the constraint set $\mathcal{C}$. Furthermore, with $T\approx\sqrt{n}$ (batch size $B=n/T$) when $\mathbf{x}^{\dagger}\not\in\mathcal{C}$. Our algorithm exactly two gradient calls per data point, and is single epoch, and does not make any convexity assumption for privacy guarantees.
@@ 2.2 DP-FTRL : DP follow the regularized leader
DP-FTRL  [44, 34, 18] is another standard approach for optimization, where in its simplest form, the state update is exactly that of DP-SGD when $\mathcal{C}=\mathbb{R}^{p}$, except the noise in two different rounds $\mathbf{b}_{t},\mathbf{b}_{t^{\prime}}$ are not independent. Instead (in the one-dimensional setting) $(\mathbf{b}_{1},\mathbf{b}_{2},\ldots\mathbf{b}_{T})\sim\mathcal{N}(0,\left(\mathbf{C}^{\top}\mathbf{C}\right)^{-1})$, where $\mathbf{C}$ is chosen to minimize the error on prefix sums of $\mathbf{x}_{t}$. A canonical and asymptotically optimal implementation of DP-FTRL  is the binary tree mechanism of [19, 16]. While DP-FTRL has strong empirical properties comparable to DP-SGD , it is unknown whether it can achieve optimal DP-SCO error unless under additional restrictions (e.g., realizability [3]). A line of work (e.g. [13, 29, 30]) has looked at optimizing the choice of $\mathbf{C}$. These works offer constant-factor improvements over the binary tree mechanism. While these improvements are meaningful practically, we aim for asymptotic guarantees in this paper, and so we will restrict our choice of correlated noise to the binary tree mechanism for simplicity.
@@ 3 Our Algorithm
Our algorithm is a DP variant of Nesterov accelerated SGD [38, 8] using stochastic recursive gradients (SRGs): similarly to [41, 50] we maintain $\nabla_{t}$, an unbiased estimate of $\nabla F(\mathbf{x}_{t})$, and use $\nabla_{t}$ and batch $t+1$ to obtain $\nabla_{t+1}$. Our algorithm incorporates noise $\mathbf{b}_{t}$ for DP. We first give a utility analysis for a fixed value of the noise $\mathbf{b}_{t}$, and then in Section 5.2 we will specify how to set the noise to simultaneously achieve good privacy and utility. As mentioned in Section 1, we will use the binary tree mechanism to set the $\mathbf{b}_{t}$’s.
It is worth mentioning that all three aspects of our algorithm are crucial: a) SRGs allow ensuring that the $\ell_{2}$-sensitivity of the gradient estimate at time $t$ is a factor of $1/t$, lower than that of the direct minibatch gradient ($\frac{1}{B}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t};d)$), b) Noise addition via binary tree mechanism allows the standard deviation of the noise $\mathbf{b}_{t}$ to grow only polylogarithmically in $T$, as opposed to $\sqrt{T}$ via simple composition analysis [20], and c) Accelerated SGD allows appropriately balancing the variance introduced by the SRG estimates, and that due to the i.i.d. sampling of the minibatches from the population. In fact in Section 4.3 we argue that it is necessary to have an accelerated SGD algorithm to achieve a single epoch optimal DP-SCO algorithm with batch size $B=\omega(1)$, if we were to use the SRGs as the gradient approximators.
1:Input: Steps: $T$, batches: $\{\mathcal{B}_{t}\}_{0\leq t<T}$ of size $B$, learning rates: $\{\eta_{t}\}_{0\leq t<T}$, scaling $\beta$, interpolation params: $\{\tau_{t}\}_{t\in[T]}$, $\mathcal{C}:\left\|\mathcal{C}\right\|_{2}\leq R$.
2:$\mathbf{x}_{0},\mathbf{z}_{0}\leftarrow 0$.
3:for $t\in\{0,\ldots,T-1\}$ do
4:     $\Delta_{t}\leftarrow\frac{1}{B}\sum_{d\in\mathcal{B}_{t}}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)]$
5:     $\nabla_{t}\leftarrow\frac{\sum_{i\leq t}\Delta_{i}}{\eta_{t}}$ (equivalently, $\nabla_{t}\leftarrow\frac{\Delta_{t}}{\eta_{t}}+\frac{\eta_{t-1}}{\eta_{t}}\cdot\nabla_{t-1}$)
6:
7:     $\mathbf{z}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{z}_{t}-\frac{\eta_{t}}{\beta}\nabla_{t}+\mathbf{b}_{t}\right)$, where $\mathbf{b}_{t}$ is the noise added to ensure privacy.
8:     $\mathbf{y}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}\right)$.
9:     $\mathbf{x}_{t+1}\leftarrow\left(1-\tau_{t+1}\right)\mathbf{y}_{t+1}+\tau_{t+1}\mathbf{z}_{t+1}$.
10:end for
@@
4 Overview of Analysis of Algorithm 1
In this section we provide a high-level overview of the analysis of Algorithm 1.
4.1 Utility Analysis
First, ignoring privacy concerns we provide a excess population risk guarantee of Algorithm 1 for any choice of the privacy noise random variables $\{\mathbf{b}_{t}\}$’s. We only assume that the variance (over the sampling of the data samples $d\sim\mathcal{D}$) in any individual term in the SRGs (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$) is bounded by $S^{2}$.
Eventually in 4.2 we will combine Theorem 4.1 below with the privacy analysis of Algorithm 1 to arrive at the final DP-SCO guarantee, with the appropriate choice of the batch gradient steps, and the batch size. In particular, there we instantiate the noise random variables $\{\mathbf{b}_{t}\}$’s with the (by now standard) binary tree mechanism [16, 19].
Fix some setting of the $\mathbf{b}_{t}$ such that $\left\|\mathbf{b}_{t}\right\|_{2}\leq b_{max}$ for all $t$. For an appropriate setting of $\eta_{t},\tau_{t}$ and $\beta\geq M$, suppose $\mathbb{E}[\left\|\Delta_{t}-\mathbb{E}[\Delta_{t}]\right\|_{2}^{2}]\leq S^{2}$. Then, with high probability over the randomness of the batches $\mathcal{B}$, the following is true for Algorithm 1:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=$
$\widetilde{O}\left(\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{SR}{T^{3/2}}+\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}R}{T^{2}}+\frac{\beta}{T^{2}}R^{2}\right).$
Interpretation of Theorem 4.1: To understand the bound in Theorem 4.1, we can view it as three groups of terms. The first group is $\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{SR}{T^{3/2}}$, which is the error due to stochasticity of the gradients. The second is $\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}R}{T^{2}}$, which is the error due to noise added for privacy. The last is $\frac{\beta}{T^{2}}R^{2}$, which is the optimization error, i.e. the error of the algorithm without stochasticity or DP noise. Notice that the optimization error matches that of the non-stochastic part of non-private Nesterov’s acceleration [36]. Without privacy, the stochastic part of Nesterov’s acceleration depends on $\frac{V}{\beta\sqrt{T}}$ [36], where $V$ bounds stochasticity in evaluating the population gradient:
@@ 4.1 Utility Analysis
$\mathbb{E}_{\mathcal{B}_{t}}\left[\left\|\frac{1}{|\mathcal{B}_{t}|}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x};d)-\nabla F(\mathbf{x};d)\right\|_{2}^{2}\right]\leq V^{2}.$
Morally, the parameter $S$ in Theorem 4.1 can be compared to $V$. However, $S$ encodes the additional variance due to stochastic recursive gradients. Modulo the difference between $S$ and $V$, Theorem 4.1 cleanly decomposes the non-private stochastic Nesterov’s utility and the additional error due to the explicit noise introduced due to DP.
Proof overview: We define the following potential. Let $\eta_{0:t}=\sum\limits_{i=0}^{t}\eta_{i}$, then
Independent batch gradients: In Theorem 4.1 we analyzed Algorithm 1 where the $\nabla_{t}$ are computed using SRG. The analysis can easily be extended to the standard setting where each $\nabla_{t}$ is an independent batch gradient. We give this analysis in Section 5.5. To the best of our knowledge, a direct potential-based analysis of stochastic accelerated gradient descent has not appeared in the literature before besides [47]. Our analysis is simpler than that of [47], so it may be of independent interest.
4.2 Privacy analysis and DP-SCO bounds
Bounding gradient differences: We next need to bound the norms of the gradient differences for a single example $\Delta_{t}(d):=\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$. This gives us a bound on the variance $S^{2}$ in the loss bound of Theorem 4.1, and a bound on the sensitivity of the $\Delta_{t}$, which will let us bound the noise $\mathbf{b}_{t}$ needed for privacy.
By smoothness and Lipschitzness, for any $d$ we have:
$\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$ $=\eta_{t}(\nabla f(\mathbf{x}_{t},d)-\nabla f(\mathbf{x}_{t-1},d))+(\eta_{t}-\eta_{t-1})\nabla f(\mathbf{x}_{t-1},d)$
$\leq\eta_{t}M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}+(\eta_{t}-\eta_{t-1})L.$
We will choose $\eta_{t}=t+1$, so the second term is just $L$. The main challenge is then to bound the first term. We will choose $\tau_{t}\approx 1/t$ to be vanishing, so
$\mathbf{x}_{t}=(1-\tau_{t})\mathbf{y}_{t}+\tau_{t}\mathbf{z}_{t}\approx\mathbf{y}_{t},$
and in turn
$\eta_{t}M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}\approx\eta_{t}M\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|_{2}\leq\left\|\frac{\eta_{t}M}{\beta}\cdot\nabla_{t}+\frac{M}{\beta}\mathbf{b}_{t}\right\|_{2}.$
The term $\frac{M}{\beta}\mathbf{b}_{t}$ ends up having norm at most $L$ under the right parameter choices, so the main challenge is to bound $\frac{\eta_{t}M}{\beta}\cdot\nabla_{t}$. Our approach for bounding this term differs based on whether or not we assume $\mathbf{x}^{\dagger}\in\mathcal{C}$. If $\mathbf{x}^{\dagger}\in\mathcal{C}$, $\mathbf{x}^{*}=\mathbf{x}^{\dagger}$ and by smoothness we have
$\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}\leq\sqrt{2M(F(\mathbf{x}_{t})-F(\mathbf{x}^{*}))}.$
Then since $\nabla_{t}$ is a stochastic estimate of $\nabla F(\mathbf{x}_{t})$, by Theorem 4.1 and a concentration inequality on $\nabla_{t}$, the fact that $F(\mathbf{x}_{t})-F(\mathbf{x}^{*})$ implies a bound on $\left\|\nabla_{t}\right\|_{2}$ that is also decreasing. In turn, while $\eta_{t}$ is increasing, we are able to show $\frac{\eta_{t}M}{\beta}\cdot\nabla_{t}$ remains constant even with $\beta=O(M)$ (Lemma 5.5). A subtle challenge is that in order to obtain a bound on $\left\|\nabla_{t}\right\|_{2}$ that is decreasing, we actually employ a tighter version of Theorem 4.1 that uses the exact norms $\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}$ rather than the Lipschitz constant $L$ in the final bound. In turn, there is a circular dependence between the bound of Theorem 4.1 and the norm of the gradient differences. Because of the circular dependence, bounding both simultaneously requires a proof by induction with careful choice of parameters (Lemma 5.4).
Without assuming $\mathbf{x}^{\dagger}\in\mathcal{C}$, it’s no longer the case that $\nabla_{t}$ is necessarily decreasing over time (e.g. in the case of linear losses, $\nabla_{t}$ will remain constant). In this case, we more directly bound $\frac{\eta_{t}M}{\beta}\cdot\nabla_{t}$ by choosing $\beta\geq TM\geq\eta_{t}M$, which ensures the gradient differences still have constant sensitivity (Lemma 5.9). The fact that we have to increase $\beta$ (i.e., slow down learning) means we need to increase $T$ to compensate in utility, leading to the increased iteration complexity $T=\sqrt{n}$ instead of $T\approx n^{1/4}$.
Privatizing the gradient differences: Once we have a bound on $\Delta_{t}(d)$, we now need to bound $\max_{t}\left\|\mathbf{b}_{t}\right\|_{2}$, i.e. the maximum noise added to any gradient estimate $\nabla_{t}$. Since each $\nabla_{t}$ is a (rescaling of) prefix sums of $\Delta_{t}$, we do this using the binary tree mechanism of [19], which adds noise to the prefix sums with variance scaling as ${\rm polylog}\left(T\right)$ rather than $T$, what we would get by independently noising each $\Delta_{t}$. We can use existing results on the binary tree mechanism (for convenience, reproven as Lemma 5.6) to obtain bounds on the maximum noise $\max_{t}\left\|\mathbf{b}_{t}\right\|_{2}$ needed to achieve $(\varepsilon,\delta)$-DP.
Finally, it is straightforward to get our main results by plugging in the bounds on $\max_{t}\left\|\mathbf{b}_{t}\right\|_{2}$ and $\Delta_{t}(d)$ into Theorem 4.1:
Theorem 4.2 (Simplified version of Theorems 5.8 and 5.12).
For appropriate settings of $\beta,T$, Algorithm 1 is $(\varepsilon,\delta)$-DP and satisfies:
$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})=O\left(LR\cdot{\rm polylog}\left(n,1/\delta\right)\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{p}}{\varepsilon n}\right)\right).$
If we assume $\mathbf{x}^{\dagger}\in\mathcal{C}$, then the parameter settings are such that $T=\widetilde{\Theta}(n^{1/4})$ and $BT=n$, i.e. only a single epoch is required. Without this assumption, the parameter settings are such that we can allow any $T\geq\sqrt{n}$ and $BT=n$, i.e. still only a single epoch is required and the batch size can be as large as $\sqrt{n}$.
4.3 Unaccelerated Variant
A natural question is whether an unaccelerated variant of Accelerated-DP-SRGD can also achieve optimal DP-SCO rates with batch size $\omega(1)$. We give a sketch that the unaccelerated version requires batch size $O(1)$. Suppose we use $\Delta_{t}=\eta_{t}\nabla f(\mathbf{x}_{t},\mathcal{B}_{t})-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},\mathcal{B}_{t})$ as our gradient differences and $\nabla_{t}=\frac{\sum_{i\leq t}\Delta_{i}}{\eta_{t}}$ as our stochastic recursive gradients. For simplicity we will consider using SGD with a fixed learning rate $1/\beta$ as the underlying optimization algorithm, assume $L=R=M=1$ and assume $\eta_{t}$ grows at most polynomially, which implies $\sum_{i\leq t}\eta_{i}^{2}\approx t\eta_{t}^{2}$. Similarly to our sensitivity bound for the accelerated algorithm (Lemma 5.9), we can show the sensitivity of each $\Delta_{t}$ is roughly $\eta_{t}/\beta$. Suppose we use $1/\beta\approx\sqrt{B/T}=B/\sqrt{n}$, which is standard for interpolation between full-batch GD and SGD with $B=1$ (see e.g. Section 6 of [12]). We compute the vector variance $\mathbf{Var}\left(\mathbf{v}\right):=\mathbb{E}[\left\|\mathbf{v}-\mathbb{E}[\mathbf{v}]\right\|_{2}^{2}]$ of $\nabla_{t}$:
$\mathbf{Var}\left(\nabla_{t}\right)=\mathbf{Var}\left(\frac{\sum_{i\leq t}\Delta_{t}}{\eta_{t}}\right)=\frac{\sum_{i\leq t}\mathbf{Var}\left(\Delta_{t}\right)}{\eta_{t}^{2}}=\frac{\sum_{i\leq t}\eta_{i}^{2}}{\eta_{t}^{2}\beta^{2}}=\frac{t}{B\beta^{2}}=\frac{tB}{n}.$